Circular No. 8843

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2007ca–2007ck Numerous supernovae have been discovered on unfiltered CCD frames:

2007cd by K. Itagaki (cf. IAUC 8817; via S. Nakano); 2007ce by R. Quimby (cf. IAUC 8779); 2007cf and 2007ci by T. Puckett (cf. IAUC 8834; 2007cf with W. McDonald, 2007ci with T. Crowley); 2007ch by L. A. G. Monard (cf. IAUC 8814); 2007cj by T.-H. Liou, T.-w. Chen et al. (cf. IAUC 8834); 2007ck by T. Boles (cf. IAUC 8789); and the rest by W. Li via LOSS/KAIT (cf. IAUC 8829; 2007ca with R. Prasad; 2007cb, 2007cc, and 2007cg with P. Thrasher). Discovery observations:

SN  2007 UT  α_{2000}  δ_{2000}  Mag.  Offset
2007ca Apr. 25.34 13^{h}31^{m}05.^{s}81 −15^{o}06′06′′.6 16.7 25′′.1 E, 1′′.7 S
2007cb Apr. 27.35 13 58 17.20 −23 22 20.7 16.8 0′′.9 E, 3′′.5 S
2007cc Apr. 27.35 14 08 41.98 −21 35 48.4 15.9 1′′.2 E, 2′′.7 N
2007cd Apr. 27.61 13 29 26.48 +11 00 44.8 17.5 9′′ E, 16′′ N
2007ce May 4.14 12 10 17.96 +48 43 31.5 17.4  —
2007cf May 8.31 15 23 07.66 + 8 31 45.5 17.9 3′′.3 E, 3′′.1 N
2007cg May 11.34 13 25 33.58 −24 39 08.1 17.7 3′′.1 E, 0′′.7 N
2007ch May 11.92 15 49 47.82 −29 23 13.7 17.2 23′′ W, 1′′ S
2007ci May 15.12 11 45 45.85 +19 46 13.9 17.5 3′′.8 W, 12′′.4 S
2007cj May 3.54 9 59 55.76 −29 37 03.3 15.9 1′′.1 E, 0′′.3 S
2007ck May 19.04 18 23 05.59 +29 54 01.0 18.5 17′′.9 E, 4′′.2 S

Additional information is available on Electronic Telegrams: 2007ca in MCG –02-34-61, CBETs 945, 947 (type-Ia supernova, discovered near maximum); 2007cb in ESO 510-G31, CBETs 949, 955 (type-Ia, discovered shortly past maximum); 2007cc in ESO 578-G26, CBETs 949, 956 (type-Ia, discovered shortly past maximum); 2007cd in NGC 5174, CBET 950; 2007ce, CBET 953 (type-Ic); 2007cf in MCG +02-39-21, CBETs 954, 958 (type-Ia, discovered several days past maximum); 2007cg in ESO 508-G75, CBETs 960, 962 (type-Ia, discovered before maximum); 2007ch in NGC 6000, CBETs 961, 963 (type-II, discovered a couple of weeks past explosion); 2007ci in NGC 3873, CBETs 966, 967 (subluminous type-Ia, discovered a couple of days before maximum); 2007cj in IC 2531, CBETs 968, 969 (type-Ia, discovered a week or two past maximum); 2007ck in MCG +05-43-16, CBETs 970, 971 (type-II, discovered a few weeks past explosion). A. Klotz writes that TAROT images of NGC 3873 taken at the Haute Provence Observatory on May 15.06 UT show nothing at the position of 2007ci (limiting mag 18.5).

2007 June 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green